Astronauts use a centrifuge to simulate acceleration of a rocket launch. The centrifuge takes 20.0 seconds to speed up from rest to its top speed of 1 rotation every 1.10 seconds. the astronaut is strapped into a seat 6.30 m from the axis. what is the astronauts tangential acceleration during the first 20.0 seconds? How many g's of acceleration does the astronaut experience when the device is rotating at top speed? Each 9.80 m/s2 of acceleration is 1 g
Answer #1

initial angular speed = wo = 0

final angular speed w1 = 2pi/T = 2pi/1.1 = 5.712 rad/s


initial tangential speed v1 = 0

radius = r = 6.3 m


(a)

tangential speed = v2 = r*w2 = 6.3*5.712 = 35.986 m/s <<-----------answer


(b)


tangential acceleration = (v2-v1)/t = 1.8 m/s^2 <<-----------answer

radial acceleration at top speed = v2^2/r = 205.6 m/s^2 <<-----------answer

a = (205.6/9.8) = 20.98 g <<-----------answer